# ADA-DPM: A Neural Descriptors-based Adaptive Noise Filtering Strategy for SLAM

### v2

@@ 4.1 Implementation Details
4.2 Compared with Others
Figure 7: The memory consumption comparison with the map representation of KISS-ICP.
Figure 8: Visualization results of trajectory estimated and map reconstruction compared with the map representation of KISS-ICP.
To validate the effectiveness of our method, we conducted experiments on several datasets (SemanticKITTI, KITTI-360, and MulRan) and compared our results with seven state-of-the-art odometry and SLAM methods: KISSICP[51], LeGO-LOAM[24] and its follow-up work SC-LeGO-LOAM[47], Mulls[48], CT-ICP[49], DeepPointMap[14], and the advanced point cloud registration method GeoTransformer[50].
Trajectory Estimated: As shown in Figure 6, ADA-DPM achieves high-precision and globally consistent map reconstruction on all selected datasets. In the figure, the red line represents the ground-truth trajectory, the blue line represents the trajectory predicted by ADA-DPM, and the red pentagram marks the starting point of the trajectory. In particular, for long-range localization and mapping tasks like KITTI-360 (where KITTI-360: 09 contains point clouds over 10,000 meters), ADA-DPM still demonstrates favorable performance.
The trajectory estimation comparison results of ADA-DPM with other algorithms on 12 test sequences are shown in Table 1. As seen from Table 1, ADA-DPM achieved the best performance on 5 sequences and the second-best performance on 3 sequences (due to CT-ICP being able to run only based on its own provided data, thus it did not participate in the testing of all sequences).
@@ 4.2 Compared with Others
Map Reconstruction: To evaluate the map representation efficiency of ADA-DPM during large-scale map reconstruction, we compared the localized storage data sizes of ADA-DPM and KISS-ICP on two large-scale sequences (SemanticKITTI: 08 and KITTI-360: 09; all data are stored in float32 data type). The results are shown in Figure 7. Benefiting from ADA-DPM’s adoption of deep learning-based sparse point cloud feature descriptors (we used the same map representation as DeepPointMap), ADA-DPM reduced memory usage by 60% compared to KISS-ICP on SemanticKITTI:08 and approximately 50% on KITTI-360:09.
Furthermore, to further evaluate the performance of ADA-DPM in SLAM task, we visualized the trajectories predicted by ADA-DPM and KISS-ICP, as well as the reconstructed maps, on SemanticKITTI:08 and KITTI-360:09. In Figure 8, the first column shows the ground-truth map, i.e., the global map reconstructed from the original point cloud using ground-truth poses; the second and third columns present the reconstructed maps by ADA-DPM and KISS-ICP, respectively; the last column displays the trajectories predicted by ADA-DPM and KISS-ICP. As shown in the figure, compared to KISS-ICP, ADA-DPM constructs more accurate global maps using less memory and achieves higher localization accuracy. Especially for long-range sequences like KITTI-360:09, ADA-DPM realizes more precise globally consistent localization and mapping compared to KISS-ICP.
Table 2: Trajectory estimated results ($APE\downarrow$) on the validation sequences.The bold indicates the best result
@@
A 2.26$\setminus$3.29 7.96$\setminus$10.21 6.02$\setminus$7.01
B $\surd$ 2.01$\setminus$2.72 8.02$\setminus$10.82 5.21$\setminus$7.44
C $\surd$ $\surd$ 2.64$\setminus$2.84 5.16$\setminus$5.62 5.56$\setminus$6.94
D $\surd$ $\surd$ $\surd$ 0.91$\setminus$1.67 5.26$\setminus$5.28 3.26$\setminus$4.32
4.3 Ablation experiments
@@
[…]
Compared Method A with Method B, it can be observed that for urban scenes such as sequences 00 and 02 (scenes with a high density of dynamic objects like vehicles and pedestrians), the addition of the Dynamic Segmentation Head effectively filters out interference from dynamic objects, leading to a reduction in APE. When comparing Method B with Method C, although the APE on sequences 00 and 02 slightly increases after incorporating the Importance Scoring Head, under conditions with added random noise, Method C exhibits a significantly lower APE than Method B on sequence 01, while the difference between the two methods is minimal on sequences 00 and 02. The reason may lie in the fact that for scenes with frequent loop closures (e.g., sequences 00 and 02), loop closure detection can mitigate cumulative errors caused by noise points to some extent; however, for scenes with fewer loop closures (e.g., sequence 01), it is challenging to eliminate cumulative errors through loop closure detection alone. Method C, by leveraging the Importance Scoring Head to assign contribution scores to each feature point pair, can mitigate the impact of noise points to a certain degree, with more pronounced effects particularly in scenes with fewer loop closures. Finally, comparing Method D with Method C, although Method D shows a slight increase in APE on sequence 01, the added GLI-GCN module further enhances the mapping accuracy in loop-closure regions for sequences 00 and 01.
@@
Finally, to further validate the effectiveness of the GLI-GCN module in regions with loop closures, we visualize the mapping results of the full ADA-DPM and the ADA-DPM with the GLI-GCN module removed, as shown in Figure 9. In Figure 9, positions A, B, C, and D represent four loop-closure locations, where A1, B1, C1, and D1 denote the mapping results of ADA-DPM, and A2, B2, C2, and D2 denote the mapping results of ADA-DPM without the GLI-GCN module. As shown in Figure 9, the maps constructed by ADA-DPM at each loop-closure position exhibit consistent alignment. This successful alignment is attributed to the GLI-GCN module, which utilizes neighborhood receptive fields of varying sizes to generate local feature descriptors fused with different receptive fields. This design mitigates the issue of reduced similarity of neighborhood structures in overlapping regions caused by a single receptive field, further enhancing localization and mapping accuracy in overlapping areas.

### v3

@@ 4.2 Compared with Others
Figure 7: Visualization results of trajectory estimation and map reconstruction compared with the map representation of KISS-ICP.
Figure 8: Comparison of memory consumption and runtime for the map representation of KISS-ICP.
The trajectory estimation comparison results of ADA-DPM with other algorithms on 12 test sequences are shown in Table 2. As seen from Table 1, ADA-DPM achieved the best performance on 5 sequences and the second-best performance on 4 sequences (due to CT-ICP being able to run only based on its own provided data, thus it did not participate in the testing of all sequences).
@@
Map Reconstruction: To further evaluate the performance of ADA-DPM in map reconstruction task, we visualized the trajectories predicted by ADA-DPM and KISS-ICP, as well as the reconstructed maps, on SemanticKITTI:08 and KITTI-360:09. In Figure 7, the first column shows the ground-truth map, i.e., the global map reconstructed from the original point cloud using ground-truth poses; the second and third columns present the reconstructed maps by ADA-DPM and KISS-ICP, respectively; the last column displays the trajectories predicted by ADA-DPM and KISS-ICP. As shown in the figure, compared to KISS-ICP, ADA-DPM constructs more accurate global maps using less memory and achieves higher localization accuracy. Especially for long-range sequences like KITTI-360:09, ADA-DPM realizes more precise globally consistent localization and mapping compared to KISS-ICP.
Figure 9: Visualization of feature point sampling.
Table 3: Trajectory estimation results ($APE\downarrow$) on the validation sequences.The bold indicates the best result
@@
[…]
Furthermore, to evaluate the map representation efficiency of ADA-DPM during large-scale map reconstruction, we compared the localized storage data sizes of ADA-DPM and KISS-ICP on two large-scale sequences (SemanticKITTI: 08 and KITTI-360: 09; all data are stored in float32 data type). The results are shown in Figure 8(a). Benefiting from ADA-DPM’s adoption of deep learning-based sparse point cloud feature descriptors (we used the same map representation as DeepPointMap), ADA-DPM reduced memory usage by 60% compared to KISS-ICP on SemanticKITTI:08 and approximately 50% on KITTI-360:09. The runtime of ADA-DPM and KISS-ICP are shown in Figure 8(b). ADA-DPM processes a single frame in approximately 0.150.50.15–0.5 s, while KISS-ICP processes a frame in 0.010.050.01–0.05 s. Compared with KISS-ICP, ADA-DPM has no advantage in running speed. However, ADA-DPM performs better in global map reconstruction and trajectory estimation, and it also has a significant advantage in map memory usage.
@@
0-10 % 1.67 5.28 4.32 3.76 $19\%$
0-20 % 2.03 5.75 4.04 3.94 $25\%$
0-30 % 2.16 5.98 4.53 4.22 $34\%$
0-40 % 1.72 5.79 4.80 4.10 $30\%$
0-50 % 2.26 5.63 4.64 4.18 $33\%$
4.3 Ablation experiments
As shown in Table 3, we separately removed the GLI-GCN module, Importance Scoring Head, and Dynamic Segmentation Head to assess their importance. To validate the robustness of the Importance Scoring Head to noise points, we compared test results under noise and noise-free conditions. We simulated noise conditions by randomly adding noise points sampled from the normal distribution $\mathcal{N}(0,0.1^{2})$ to 0%–10% of the point cloud, with noise clipped to the range [0.2,0.2]m[-0.2,0.2]m on each axis. In Table 3: Method D refers to ADA-DPM; Method C is ADA-DPM with the GLI-GCN module removed; Method B is Method C with the Importance Scoring Head removed; Method A is Method B with the Dynamic Segmentation Head removed.
Compared Method A with Method B, it can be observed that for urban scenes such as sequences 00 and 02, which contain many dynamic objects (e.g., cars and pedestrians), adding the Dynamic Segmentation Head leads to a decrease in APE. We visualized the feature points of Method A and Method B to qualitatively analyze the effect of the Dynamic Segmentation Head. As shown in Figure 9, after adding the Dynamic Segmentation Head, only a small number of the feature points sampled by Method B correspond to dynamic objects, demonstrating that the Dynamic Segmentation Head can partially filter out the interference from dynamic objects.
When compared Method B with Method C, although the APE on sequences 00 and 02 slightly increases after incorporating the Importance Scoring Head, under conditions with added random noise, Method C exhibits a significantly lower APE than Method B on sequence 01, while the difference between the two methods is minimal on sequences 00 and 02. The reason may lie in the fact that for scenes with frequent loop closures (e.g., sequences 00 and 02), loop closure detection can mitigate cumulative errors caused by noise points to some extent; however, for scenes with fewer loop closures (e.g., sequence 01), it is challenging to eliminate cumulative errors through loop closure detection alone. Method C, by leveraging the Importance Scoring Head to assign contribution scores to each feature point pair, can mitigate the impact of noise points to a certain degree, with more pronounced effects particularly in scenes with fewer loop closures. Finally, compared Method D with Method C, although Method D shows a slight increase in APE on sequence 01, the added GLI-GCN module further enhances the mapping accuracy in loop-closure regions for sequences 00 and 01.
@@ 4.3 Ablation experiments
Finally, to further validate the effectiveness of the GLI-GCN module in regions with loop closures, we visualize the mapping results of the full ADA-DPM and the ADA-DPM with the GLI-GCN module removed, as shown in Figure 10. In Figure 10, positions A, B, C, and D represent four loop-closure locations, where A1, B1, C1, and D1 denote the mapping results of ADA-DPM, and A2, B2, C2, and D2 denote the mapping results of ADA-DPM without GLI-GCN module. As shown in Figure 10, the maps constructed by ADA-DPM at each loop-closure position exhibit consistent alignment. This successful alignment can be attributed to GLI-GCN, which utilizes different neighborhoods to perceive receptive fields of varying sizes. By fusing and emphasizing shallow features that have high correlation with deep feature points, it generates local feature descriptors with richer geometric characteristics. This alleviates the issue of reduced neighborhood structure similarity in overlapping regions caused by a single receptive field, thereby further enhancing localization and mapping accuracy in the overlapping areas.
@@ 5 Conclusions
Furthermore, compared with KISS-ICP, ADA-DPM has no advantage in terms of speed. On one hand, deep learning-based methods generally require more GPU computational power. In future work, PointNeXt could be replaced with a more lightweight network.